OpenStudy (anonymous):

Need help with a math question !! EASY 10 POINTS ANYONE WILLING TO HELP :[ ?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Let's hear it!

OpenStudy (anonymous):

3. Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (2 points)

OpenStudy (anonymous):

No. of quarters—q: 25q + 10(46 - q) = 700 5q + 2(46 - q) = 140 5q + 92 - 2q = 140 3q = 48 q = 16 No. of dimes: = 46 - 16 = 30 Answer: 16 quarters & 30 dimes

OpenStudy (anonymous):

OMGG !! HOW DID YOU DO IT? can you like explain it ?

OpenStudy (anonymous):

You did it so fast , ive spent over 3 hours trying to figure it out .

OpenStudy (anonymous):

like break it down so i can understand it , cause if i just put the answer its not gonna help me on my test lol.

OpenStudy (anonymous):

i did this math like years ago